# Software Fairness Testing in Practice

### v1

@@ V-D Threats to Validity
Finally, in this study, we used GPT-4 Omni through prompt engineering to assist with open-coding participant responses, yet this approach introduces specific limitations. The effectiveness of prompt engineering depends heavily on clarity and precision in crafting prompts, which may risk introducing bias in the qualitative analysis process. Although the few-shot examples and a defined task persona helped direct the model, the inherent variability of generative AI models presents challenges for the consistency and reproducibility of results. We mitigated this issue by conducting a 20% accuracy check to verify alignment between the codes generated by the model and the original transcripts, ensuring that any potential discrepancies were identified early. Additionally, we continued to validate this alignment manually throughout the remaining stages of the qualitative analysis process. This approach allowed us to maintain consistency and accuracy in our findings while leveraging the model’s efficiency for initial coding.
@@ VII Data Availability
Anonymized interview quotations can be accessed via this link: https://figshare.com/s/d6676bece29428ee7a25. Quotations containing direct references to participants’ names, colleagues, specific projects, or companies were excluded to ensure privacy. Consequently, some participants may not have associated quotations in the spreadsheet. Additionally, certain quotations may appear awkwardly phrased due to direct translations from participants’ native languages.

### v2

@@ V-D Threats to Validity
Finally, in this study, we used GPT-4 Omni through prompt engineering to assist with open-coding, yet this approach introduces some limitations. The effectiveness of prompt engineering depends on clarity and precision in crafting prompts, which may risk introducing bias in the qualitative analysis process. Although the few-shot examples and a defined task persona helped direct the model, the inherent variability of generative AI models presents challenges for the consistency and reproducibility of results. We mitigated this issue by conducting a manual accuracy check to verify alignment between the codes generated by the model and the original transcripts. This approach allowed us to maintain consistency and accuracy in our findings while leveraging the model’s efficiency for initial coding.
@@ VII Data Availability
Anonymized interview quotations are available at https://figshare.com/s/d6676bece29428ee7a25. Quotes with identifying details were excluded for privacy, so some participants may not have associated quotations in the spreadsheet. Finally, some phrasing may reflect direct translations from other languages.
Acknowledgments
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), Discovery Grant RGPIN-2024-06260, and by Alberta Innovates through the Advance Program, Project Number 24506125.